OpenStudy (anonymous):

what is the value of x if the ordered pair is (10,5) and (x,9) has a slope of -2?

OpenStudy (dls):

\[\LARGE m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (dls):

\[\LARGE \frac{9-5}{x-10}=-2\]

OpenStudy (dls):

Solve for x.

OpenStudy (anonymous):

Thanks

OpenStudy (dls):

x=?

OpenStudy (anonymous):

its 12 hannah just in case you wanted the answer instead like most of us do...

OpenStudy (anonymous):

no it's not because that's not one of the answers Anders

OpenStudy (anonymous):

whats the answer then cuz I really thought that was it.

OpenStudy (anonymous):

it's -2